### v1

# Phonon, Infrared and Raman Spectra of LiGa5O8 from Density Functional Perturbation Theory

Abstract
LiGa5O8 with a cubic spinel type structure was recently reported to be a ultra-wide-band-gap semiconductor with unintentional p-type conduction. While the origin of p-type doping is still unclear, the fundamental properties of this material are of interest. Here we present a first-principles study of the phonons using density functional perturbation theory. The phonon band structures show no unstable modes verifying the stability of the structure. We focus mainly on the phonons at the Brillouin zone center for which a full symmetry analysis is presented. We present the dielectric function contributions from the infrared active modes as well as the Raman spectra and their polarization dependence. The phonon density of states integrated over the Brillouin zone is also presented as this may related to disordered Raman spectra.
I INTRODUCTION
LiGa5O8 was recently reported to be a p-type ultra-wide-band-gap (UWBG) semiconductor with a gap of 5.34 eV and cubic spinel structure [1]. This is an unusual finding as native p-type doping in UWBG oxides is rare and p-type doping is often hindered by small polaron formation. The experimental gap [1] extracted from a Tauc plot is slightly smaller than the calculated gap of 5.72 eV [2] using the quasiparticle self-consistent (QS) method, which is a many-body perturbation theory (MBPT) method in which the self-energy is calculated from the one particle Green’s function and the screened Coulomb interaction including electron-hole interactions in the screening. The optical band gap obtained from the Bethe Salpeter Equation (BSE) in this same work [2] indicates a large exciton binding energy with lowest direct exciton gap of 5.5 eV and also showed the existence of a 0.1 eV lower indirect gap. The band gap is thus fairly well established. However, the origin of the p-type doping is not. Native defect calculations [3] showed all plausible acceptors to have high binding energies exceeding 0.75 eV. Furthermore, compensating shallow donors like GaLi antisites or deep donor oxygen vacancies were found to pin the equilibrium Fermi level deep in the gap thus predicting insulating behavior. Similar findings were found with evidence of small polaron formation for native defects and some candidate acceptor impurities by Lyons [4]. Insulating behavior has meanwhile been found upon annealing in O2 and the stoichiometry conditions under which p-type behavior occurs have been more clearly established [5] to be somewhat O-poor, Li-poor and Ga-rich. In spite of the lack of understanding of the origin of p-type doping rectification behavior at pn-heterojunçtions between n-type Ga2O3 and p-type LiGa5O8 has been reported [6]. LiGa5O8 has previously received some interest as a host for phosphorescent dopants like Cr [7, 8, 9]. The spinel structure of this compound was reported by Joubert et al.[10]. While the present paper does not address directly the question of p-type doping, the renewed interest in this material motivates a broader study of its fundamental properties. Here we present a first-principles study of the phonons with the aim to assist on-going characterization efforts on this new material. First, the phonon band structure calculations are important to verify the mechanical stability of the material in this structure. Second, Raman spectra are often used to characterize the crystalline quality of a material. Third, electron-phonon coupling is responsible for finite temperature gap renormalization and for limitations to the mobility. We here focus on the phonons at the Brillouin zone center and present a symmetry analysis of the vibrational modes of the 56 atom unit cell and report the calculated Infrared and Raman spectra as well as related Born effective charges and static and high frequency dielectric constants.
II COMPUTATIONAL METHOD
All the calculation presented here were carried out using the Projector Augmented Wave (PAW) method [11] as implemented in the Abinit code [12, 13]. The following electron configurations were adapted as valence electrons: for Li, for Ga, and for O. We used the generalized gradient approximation (GGA) in the PBESol parameterization [14] as exchange correlation functional. We start from the structures provided by the Materials Project [15, 16, 17] and optimize the ground state structure first before proceeding with the electronic band structure calculation and subsequently studying the vibrational properties using density functional perturbation theory (DFPT) [18, 19]. DFPT relies on the theorem which allows one to calculate up to third order derivatives of the total energy using first-order corrected wave functions. This allows one to obtain second derivatives versus atomic positions (force constants), second and third derivatives with respect to long range electric fields (high-frequency dielectric constants and second order susceptibilities in the static but electronic screening only limit), mixed second derivatives versus electric field and atomic displacements (Born effective charges as needed for the LO-TO splitting). For the Raman intensities one needs the derivative of the electric susceptibility versus atomic displacements, which is a mixed third order derivative[20]. Within the PAW or ultrasoft pseudopotential methods, however, the calculation of the Raman tensor requires second order corrected wave functions [21]
The wave function cut-off used is 60 Ha and the PAW energy cut-off for the double grid was chosen to be 120 Ha. The structural relaxation was carried out with the tolerance for forces less than Ha/Bohr. For the Brillouin zone sampling of the elecronic properties, an unshifted Monkhorst-Pack grid of size was used. To obtain phonon dispersion and an accurate density of phonon states (DOS), we use the interpolation approach of [19] using analytic expressions for the long-range force constants and real space short-range force constants extracted from an initial coarse q-mesh by Fourier transformation. The final q-mesh used for the phonon DOS was with an energy cut-off of 25 Ha.
For, the calculation of microscopic dielectric properties, IR, and Raman spectra, a zone centered phonon calculation was performed. The convergence criterion on forces used for the phonon calculations is Ha/Bohr but for the higher order corrections in wavefunction the convergence criterion was chosen in terms of wavefunction residual of .
III RESULTS AND DISCUSSION
III.1 Crystal Structure
Figure 1 represents the spinel-type LiGa5O8 structure which contains four formula units. The unit cell contains four Li in the Wyckoff 4b positions, which are octahedrally coordinated, eight tetrahedral Ga (8c) sites, twelve octahedral Ga (12d) sites, and thirty-two O in two distinct Wyckoff sites O(1) in 8c positions and O(2) in 24e positions. In total, there are 56 atoms in the unit cell. The structure is cubic with space group (No.212) and point group in the Schoenflies notation or in the International Crystallography notation.
The calculated optimized parameters are compared with the experimental ones and previous calculations in Table 1. Excellent agreement is obtained with experiment for the cubic lattice constant with a slight 0.3 % overestimate and the reduced coordinates of the the Wyckoff positions agree well with the Material Project optimization to at least 3 digits. The bond lengths are reported in Table 1.
Parameters | Theory | Expt. |
Lattice constant a (Å) | 8.233 | 8.203[10] |
Wyckoff positions | ||
4d Li | (5/8, 5/8, 5/8) | (5/8, 5/8, 5/8) |
8c Ga | (0.253845, 0.253845, 0.253845) | (0.253299, 0.253299, 0.253299)[16] |
12d Ga | (1/8, 0.365375, 0.884623) | (1/8, 0.365985, 0.884015) |
8c O(1) | (0.112737, 0.612737, 0.887263) | (0.11325, 0.613253, 0.886747) |
24e O(2) | (0.873628, 0.884526, 0.380629) | (0.87316, 0.884637, 0.381276) |
Bond lengths (Å) | ||
Li-O | 2.072 | |
Gat-O(1) | 1.867 | |
Gat-O(2) | 1.903 | |
Gao-O(2) | 1.986 | |
Gao-O(2) | 1.939 | |
Gao-O(1) | 2.039 |
III.2 Group Theoretical Analysis
Since there are 56 atoms in the cell there are 168 degrees of motion or vibrational modes. Of these, three are acoustic and correspond to a translation of the crystal at . Group theory allows one to classify the remaining modes according to their irreducible representations. The character table of the group is given in Table 2. For each type of Wyckoff position we can derive a reducible representation from the displacements in the three Cartesian directions with the characters shown in the second part of the table. These can then be reduced to irreducible representations as shown in the last column. Since there are both tetrahedral Gat and O(1) in 8c positions, the 168 modes break up into modes. Of these one corresponds to the uniform translations and needs to be removed. Of the remaining 165 modes, only is infrared (IR) active because it corresponds to the components of a vector. Thus we find 21 IR active modes. These will show an LO-TO splitting. Raman modes correspond to a symmetric second rank tensor and can thus belong to irreducible representations. So there are Raman modes and the modes are silent. The Raman modes of symmetry only have off-diagonal non-zero components and are all three-fold degenerate. The Raman modes have a Raman tensor proportional to the unit matrix while the Raman modes are doubly degenerate and have only diagonal components as shown below.
Modes | E | 8 | 6 | 6 | 3 | Function |
1 | 1 | 1 | 1 | 1 | ||
1 | 1 | -1 | -1 | 1 | – | |
2 | -1 | 0 | 0 | 2 | ||
3 | 0 | -1 | 1 | -1 | ||
3 | 0 | 1 | -1 | -1 | ||
12 | 0 | -2 | 0 | 0 | ||
24 | 0 | 0 | 0 | 0 | ||
36 | 0 | -2 | 0 | 0 | ||
72 | 0 | 0 | 0 | 0 |
III.3 Electronic band structure
While the electronic band structure has been already studied with superior methods such as the QS method in [2] we here check the band structure obtained in the PBEsol functional. The electronic band structure is shown in fig. 2 and shows a slightly smaller indirect gap. The direct 2.86eV gap at and indirect gap 2.76eV found here are consistent with the GGA gaps given in that paper which are respectively 2.58 eV and 2.48 eV. That paper also used the PBEsol functional but a different all-electron method and the full-potential linearize muffin-tin orbital method. It shows that the convergence of our present PAW calculation is adequate.
III.4 Phonon dispersions and density of states
The phonon band structure is shown in fig. 3. First, we note that there are no imaginary frequency modes, which verifies the mechanical stability of the structure. We can recognize the three lowest acoustic modes followed by 165 modes. The dips near reflect the LO-TO splitting of the IR active modes. The values at itself represent the TO frequency while the limit represent the LO modes.
Next, we show the Phonon Density of states (DOS) in fig.4.
It shows the usual quadratic onset in the region of acoustic modes. The decomposition in atomic components shows that the highest modes have primarily O character. Even though Li atoms have smaller mass, their main contributions appear between 300-400 and near 500 cm-1. This indicates that the Li-O bonds have smaller force constants than some of the Ga-O ones. This is consistent with the bond-lengths in Table 1. The lower energy range up to 250 cm-1 shows primarily Ga contribution as expected because Ga is the heaviest atom in the system.
III.5 Phonons at the zone center
The numerical values of the phonons at are
presented in Table 3 and sorted according to their symmetry character. At the zone center (q=0), due to the long-range interaction between the macroscopic electric field and the Born effective charge LO-TO splitting is present for the modes. We can see that for the low frequency range the splitting is not so prominent, but in the high frequency ( cm-1) regime the splitting is significant. This reflects the stronger dipolar character of these modes, which correspond to the various bond stretch modes.
123.3 | 150.2 | 129.8 | 138.2 | 467.5 | 138.2 | 483.3 | 129.3 | 517.4 |
258.9 | 193.5 | 148.1 | 172.8 | 491.9 | 172.8 | 492.4 | 150.6 | 544.7 |
403.6 | 273.9 | 169.1 | 189.5 | 498.7 | 189.8 | 529.6 | 177.3 | 570.1 |
484.5 | 360.6 | 225.8 | 193.3 | 549.3 | 193.6 | 571.7 | 194.2 | 614.5 |
525.6 | 469.5 | 260.9 | 215.2 | 587.5 | 216.9 | 655.6 | 204.1 | 666.4 |
715.1 | 495.7 | 306.5 | 239.2 | 665.3 | 239.9 | 693.4 | 246.2 | 694.6 |
520.3 | 323.4 | 276.2 | 714.8 | 276.2 | 724.1 | 288.7 | ||
698.8 | 352.9 | 303.6 | 303.7 | 306.0 | ||||
379.7 | 320.6 | 331.0 | 327.6 | |||||
407.1 | 332.6 | 359.0 | 362.2 | |||||
499.8 | 368.5 | 370.8 | 394.0 | |||||
546.8 | 390.9 | 394.6 | 438.0 | |||||
593.0 | 397.7 | 424.4 | 455.8 | |||||
603.7 | 458.3 | 462.2 | 493.7 |
III.6 IR Spectra and Dielectric Properties
The infrared optical response is fully described by the macroscopic dielectric permittivity tensor, . It is obtained from the oscillator strengths , as
where is the purely electronic screening contribution. The oscillator strengths in turn depend on the Born effective charges and eigenmode eigen vectors,
Here, indicates the atom number in the unitcell and the Cartesian directions. For reference, the diagonal elements of the Born effective charge tensors are provided in Table 4 and the oscillator strenths for the modes in Table 5. We note that in a cubic crystal the oscillator strength matrix has only diagonal components and they are all equal so there is only one component per mode . The Born effective charges are defined by
They do have small off-diagonal elements which add up to zero when summed over all equivalent atoms of a given Wyckoff position type. The Born charges of Li, Ga and O are seen to be close to their nominal values of +1, +3 and respectively but not all diagonal components are equal. This is because the Born effective charges reflect site symmetry of the Wyckoff position rather than the global symmetry. The charge neutrality rule for each pair of Cartesian components is enforced by renormalizing the initial calculated ones. Of the 12 octahedral Ga, 4 have the value listed in the table and 8 have the value listed in the table as their . Of the 24 O(2), 8 have have the , 8 the and 8 the value listed. Thus . We also provide the calculated high frequency and static dielectric constants in Table6.
The poles of the dielectric function, which correspond to peaks in its imaginary part correspond to the TO modes, while the zeros of the real part correspond to the LO modes. The latter also show up as peaks of the loss function . From these, one can extract the absorption coefficient with the complex index of refraction and the normal incidence reflectance Using a fixed broadening parameter of 5 cm-1 for all modes, we obtain the dielectric function and related properties, displayed in fig- 5. Because some modes, notably modes 8(61-63), 11 (86-88), 16(117-119), have very small oscillator strength (and small TO-LO) splitting, they are not clearly visible and hence we see only 18 instead of the predicted 21 peaks. Note that the strongest mode is mode which corresponds to modes 69-71 when ordering all modes irrespective of symmetry according to energy as given in Table 5. As an example, we show the eigendisplacement pattern for mode in fig. 6. We can see that most of the cations have an upward motion while the O atoms have a downward motion. Thus, there is a significant dipole character to the overall motions consistent with a large IR active oscillator strength. There are both Li-O and Ga-O bond stretches involved in this mode and they are in phase for equivalent atoms.
Atom | |||
Li | 1.081846 | 1.081846 | 1.081846 |
Gat | 2.964809 | 2.964809 | 2.964809 |
Gao | 3.202848 | 3.199848 | 3.199848 |
O(1) | -2.142370 | -2.142370 | -2.142370 |
O(2) | -2.200546 | -2.287746 | -1.676342 |
Mode Number | = = |
10 - 12 | |
21 - 23 | |
27 - 29 | |
33 - 35 | |
40 - 42 | |
45 - 47 | |
55 - 57 | |
61 - 63 | |
69 - 71 | |
77 - 79 | |
86 - 88 | |
91 - 93 | |
97 - 99 | |
109 - 111 | |
112 - 114 | |
117 - 119 | |
126 - 128 | |
139 - 141 | |
145 - 147 | |
155 - 157 | |
166 - 168 |
4.107 | 10.959 |
III.7 Raman spectra
In fig.8 we show the Raman modes of and symmetry. Note that the - modes have significanty smaller intensity. Both of these are active in a scattering geometry with parallel incident and scattered light polarizations. We show the spectra for modes separately in fig.8 because these are active for perpendicular incident and
the Raman spectrum averaged over directions which is relevant for a polycrystalline sample is shown in fig.9.
In fig.10 we illustrate the eigen displacements of the highest frequency mode. We can see several bond stretches but now they are out of phase between different equivalent ones so that the net oscillator strength is zero. This mode preserves the symmetry of the crystal. Note that in this mode the Li atoms do not participate.
at 715.1 cm-1
We finally note that these represent first-order Raman scattering and assume crystal momentum conservation so that only modes at the zone center contribute. Often in samples with significant defect induced disorder, the momentum conservation is relaxed and one then finds a disorder induced Raman scattering which can have contributions from the whole Brillouin zone and therefore is closer to the integrated Phonon DOS which was presented in Sec.III.4.
IV CONCLUSION
A first-principles study was presented of the phonon related properties of LiGa5O8. The PAW calculations in the PBEsol exchange correlation functional provide good agreement with experimental data on the structural parameters and a band gap consistent with previous calculations. The absence of imaginary modes confirms the mechanical stability of the spinel structure. The phonons at the zone center were calculated and symmetry labeled. There are 21 infrared active modes, which could be identified in the calculated IR spectra and for which the LO-TO splitting was reported. There are 20 Raman active modes for perpendicular input and output polarizations, 14 modes and 6 modes which are Raman active under parallel polarizations and 8 silent modes. The dielectric IR and Raman spectra as well as supporting quantities such as the oscillator strengths, high-frequency and static dielectric constants, and the Born effective charges.
Data Availability The digital data related to the figures are provided in https://github.com/Electronic-Structure-Group/Phonons-LiGa5O8.
Acknowledgements.
This work was supported by the US. Air Force Research Office (AFOSR) under grant No. FA9550-22-1-0201. It made use of the High Performance Computing Resource in the Core Facility for Advanced Research Computing at Case Western Reserve UniversityReferences
- Zhang et al. [2023] K. Zhang, V. G. T. Vangipuram, H.-L. Huang, J. Hwang, and H. Zhao, Discovery of a Robust P-Type Ultrawide Bandgap Oxide Semiconductor: LiGa5O8, Advanced Electronic Materials n/a, 2300550 (2023).
- Lambrecht [2024] W. R. L. Lambrecht, Spinel LiGa5O8 prospects as ultra-wideband-gap semiconductor: Band structure, optical properties, and doping, Journal of Vacuum Science & Technology A 42, 022705 (2024).
- Dabsamut et al. [2024] K. Dabsamut, K. Takahashi, and W. R. L. Lambrecht, Native defects and their complexes in spinel LiGa5O8, Journal of Applied Physics 135, 235707 (2024).
- Lyons [2024] J. L. Lyons, Deep polaronic acceptors in LiGa5O8, Journal of Applied Physics 135, 165705 (2024).
- Zhang et al. [2025] K. Zhang, V. G. Thirupakuzi Vangipuram, and H. Zhao, Experimental analysis of potential origin of p-type conductivity in LiGa5O8, APL Materials 13, 041104 (2025).
- Thirupakuzi Vangipuram et al. [2025] V. G. Thirupakuzi Vangipuram, K. Zhang, D. S. Yu, L. Meng, C. Chae, Y. Xu, J. Hwang, W. Lu, and H. Zhao, Ultrawide bandgap LiGa5O8/-Ga2O3 heterojunction p–n diode, APL Electronic Devices 1, 016115 (2025).
- De Clercq et al. [2017] O. Q. De Clercq, L. I. D. J. Martin, K. Korthout, J. Kusakovskij, H. Vrielinck, and D. Poelman, Probing the local structure of the near-infrared emitting persistent phosphor LiGa5O8:Cr3+, J. Mater. Chem. C 5, 10861 (2017).
- Sousa and Carvalho [2020] O. Sousa and I. Carvalho, Theoretical study of the structural, energetic, electronic and magnetic properties of the host matrix LiGa5O8 doped with Cr3+, Journal of Solid State Chemistry 289, 121472 (2020).
- Huang et al. [2018] W. Huang, X. Gong, R. Cui, X. Li, L. Li, X. Wang, and C. Deng, Enhanced persistent luminescence of LiGa5O8:Cr3+ near-infrared phosphors by codoping Sn4+, Journal of Materials Science: Materials in Electronics 29, 10535 (2018).
- Joubert et al. [1963] J. Joubert, M. Brunel, A. Waintal, and A. Durif, Etude cristallographique du gallate de lithium et de sa solution solide avec l‘aluminate, Comptes Rendus Hebdomadaires des Seances de l’Academie des Sciences (1884 - 1965) 256, 5324–5326 (1963).
- Blöchl [1994] P. E. Blöchl, Projector augmented-wave method, Physical Review B 50, 17953 (1994).
- Gonze et al. [2020] X. Gonze, B. Amadon, G. Antonius, F. Arnardi, L. Baguet, J.-M. Beuken, J. Bieder, F. Bottin, J. Bouchet, E. Bousquet, N. Brouwer, F. Bruneval, G. Brunin, T. Cavignac, J.-B. Charraud, W. Chen, M. Côté, S. Cottenier, J. Denier, G. Geneste, P. Ghosez, M. Giantomassi, Y. Gillet, O. Gingras, D. R. Hamann, G. Hautier, X. He, N. Helbig, N. Holzwarth, Y. Jia, F. Jollet, W. Lafargue-Dit-Hauret, K. Lejaeghere, M. A. L. Marques, A. Martin, C. Martins, H. P. C. Miranda, F. Naccarato, K. Persson, G. Petretto, V. Planes, Y. Pouillon, S. Prokhorenko, F. Ricci, G.-M. Rignanese, A. H. Romero, M. M. Schmitt, M. Torrent, M. J. van Setten, B. V. Troeye, M. J. Verstraete, G. Zérah, and J. W. Zwanziger, The Abinit project: Impact, environment and recent developments, Comput. Phys. Commun. 248, 107042 (2020).
- Torrent et al. [2008] M. Torrent, F. Jollet, F. Bottin, G. Zérah, and X. Gonze, Implementation of the projector augmented-wave method in the ABINIT code: Application to the study of iron under pressure, Computational Materials Science 42, 337 (2008).
- Perdew et al. [2008] J. P. Perdew, A. Ruzsinszky, G. I. Csonka, O. A. Vydrov, G. E. Scuseria, L. A. Constantin, X. Zhou, and K. Burke, Restoring the Density-Gradient Expansion for Exchange in Solids and Surfaces, Phys. Rev. Lett. 100, 136406 (2008).
- Jain et al. [2013] A. Jain, S. P. Ong, G. Hautier, W. Chen, W. D. Richards, S. Dacek, S. Cholia, D. Gunter, D. Skinner, G. Ceder, and K. A. Persson, Commentary: The Materials Project: A materials genome approach to accelerating materials innovation, APL Materials 1, 011002 (2013).
- [16] Data retrieved from Materials Project for LiGa5O8 (mp-28146) from database version v2022.10.28.
- [17] International Crystallographic Data Base, id=33716.
- Gonze [1997] X. Gonze, First-principles responses of solids to atomic displacements and homogeneous electric fields: Implementation of a conjugate-gradient algorithm, Phys. Rev. B 55, 10337 (1997).
- Gonze and Lee [1997] X. Gonze and C. Lee, Dynamical matrices, Born effective charges, dielectric permittivity tensors, and interatomic force constants from density-functional perturbation theory, Phys. Rev. B 55, 10355 (1997).
- Veithen and Ghosez [2005] M. Veithen and P. Ghosez, Temperature dependence of the electro-optic tensor and refractive indices of from first principles, Phys. Rev. B 71, 132101 (2005).
- Miwa [2011] K. Miwa, Prediction of Raman spectra with ultrasoft pseudopotentials, Phys. Rev. B 84, 094304 (2011).